# GAPS: Guiding Dynamic Android Analysis with Static Path Synthesis

### v1

@@ 3.3. Call Graph Generation
FlowDroid (flowdroid) is a taint analysis tool that generates the app’s call graph to discover privacy and security leaks between defined sources and sinks. FlowDroid builds a call graph based on Soot (soot) and performs a flow- and context-sensitive analysis. However, FlowDroid did not model ICC in its initial version, leaving an essential contribution to call graphs undressed. This was later introduced with the integration of IccTA (iccta). Similarly, Amandroid (amandroid) is a taint analysis tool that builds an inter-procedural call graph and computes points-to information for all objects of an app component. FlowDroid and Amandroid employ data-flow analysis techniques to develop a call graph in a forward-visit approach, i.e., every branch is visited during construction. In addition, this approach has proven to be resource-intensive when integrated with the increasing complexity of apps (mudflow). GAPS aims at a more scalable approach: we build a partial inter-procedural call graph backward by starting from a target method invocation to limit the processing overhead.
DroidReach (droidreach) is the only known path reconstruction tool. Its path reconstruction techniques rely on the generation of a complete call graph using either the Androguard call graph creation tool (androguard_cg) or FlowDroid. In subsection 4.3, we compare GAPS against DroidReach and FlowDroid. In both cases, GAPS’s static analysis module proved to be faster and more effective at reaching target methods compared to its counterparts. Additionally, it provides specific app entry points, execution paths to satisfy conditional statements, and corresponding GUI events to guide automated interaction.
@@ 3.4. Hybrid analysis
Andrubis (andrubis) is a hybrid analysis tool created for detecting Android malware. Andrubis’ static analysis retrieves meta information (e.g., permissions, components, package name) and examines the Dalvik bytecode to extract a complete list of Java objects and methods. Subsequently, dynamic analysis is aided by information previously collected to interact with known entry points and measure code coverage, but it is not ultimately guided by the data collected during static analysis. AndroPyTool (andropytool) is another malware detection tool that combines static and dynamic analysis. It uses Androguard to extract essential features (e.g., package name, permissions, opcodes, API calls), FlowDroid to perform taint analysis, and samples are run inside a sandbox. However, dynamic analysis is not assisted by static analysis; instead, they are included as separate contributions in the output. GAPS aims to use static analysis information to provide path-sensitive instructions and guide automatic interaction. Evodroid (evodroid) proposes an evolutionary approach to the system testing of Android apps. It builds a Control Graph Model consisting of code segments used during GUI exercising to increase coverage. GAPS, instead, proposes a targeted approach to reach a specific method inside the app. Furthermore, Evodroid bases its static analysis on MoDisco (modisco), which requires the source code of the app to be provided, which is not always possible for most apps on the Google Play Store. GAPS can analyze any APK and DEX files as input and does not require the source code to operate. LetterBomb (letterbomb) employs a hybrid approach for automatically generating exploits for Android apps, specifically targeting Inter-Component Communication (ICC) vulnerabilities. The methodology combines a path-sensitive symbolic execution-based static analysis with software instrumentation and test oracles to identify and exploit weaknesses in the ICC interface. GAPS differs from LetterBomb in that it performs static analysis without relying on symbolic execution, which is known to be computationally expensive and often impractical for large, real-world apps. Instead, GAPS uses a lightweight static approach to reconstruct paths to a target method and then translates them into high-level instructions to guide UI exploration.
@@
GoalExplorer (goalexplorer) is a hybrid system for automated, goal-driven exploration of Android applications. Its core idea is to first statically model the app’s UI screens and the transitions between them by combining results from modified versions of FlowDroid (flowdroid) and BackStage (backstage), creating an STG. This STG is then used to guide dynamic exploration, allowing the tool to navigate to a specific target of interest, such as an Android activity, an API call, or a program statement. In subsection 4.2, we have compared GoalExplorer against GAPS and it only reached 9.69% of the target methods on average over three distinct runs.
@@ 4. Motivation
RQ1: How effective are GUI testers at reaching specific target methods?
RQ2: How effective is static analysis in reaching a method?
RQ3: How does GAPS perform compared to dynamic and static approaches?
RQ4: Can GAPS be applied to real-world apps and practical use cases?
@@ 4.1. Datasets & Setup
For the first three research questions (RQ1–RQ3), we used 56 applications from the AndroTest benchmark, for which we collected the apps’ latest version. These apps were instrumented using AndroLogs (samhi2024androlog) to measure method reachability by logging every time a method was executed. For each app, we randomly selected a consistent set of 50 target methods to be reached during testing, out of which only 34.39% of them are located inside an Activity.
@@
Static analysis experiments are performed on a CloudLab (Duplyakin+:ATC19) Ubuntu 22.04 machine with 64 GB of RAM, while dynamic ones are performed on an x86-64 emulator running Android 13 and Pixel 2 with Android 11 when an ARM architecture is required.
4.2. RQ1: GUI testers effectiveness
@@
To evaluate the effectiveness of GUI testers in reaching targeted methods within Android applications, we selected APE, recognized as one of the best-performing tools in recent studies (wang2021infrastructure; su2021benchmarking; akinotcho2024mobile), and Guardian, an LLM-based UI exploration tool that has outperformed other task-guided approaches (guardian). We also compared against GoalExplorer (goalexplorer), a hybrid analysis tool that statically builds an STG to guide the Stoat (stoat) stochastic GUI testing. The experiment was conducted using the AndroTest benchmark. Each app was executed for five minutes across three independent runs. We chose a 5-minute timeout for all competing tools, including GAPS, because we aimed to have an efficient dynamic analysis. Previous works (jordanfiveminutes; marvel) chose the same threshold. During execution, logs were collected to verify the dynamic reachability of the target methods. To execute Guardian, we tasked the LLM model to “interact with the application to maximize exploration”.
@@
Considering the set of 50 target methods for each app, over three runs, GoalExplorer reached an average of 9.69% methods, APE 12.82%, and Guardian 17.12%. These results show how the three techniques differ in usability: APE can rapidly generate GUI events, but is unable to consistently build app states that can help discover more hidden parts of the app code. GoalExplorer uses the STG to guide Stoat by flagging activities when a target API is invoked. However, it faces limitations when the target methods reside in third-party libraries or when control-flow conditions are involved, as these often require more complex interactions beyond merely reaching the target activity. Finally, Guardian takes advantage of the LLM decision-making capabilities to systematically interact with UI elements in a goal-oriented manner, enabling it to uncover deeper execution paths. Despite this, its average reachability remains low, highlighting that semantic planning alone is insufficient without deeper control-flow awareness.
@@
RQ1: GUI testers effectiveness Our evaluation highlights the limitations of dynamic interaction tools in reaching target methods. Both APE and Guardian show low success rates, with GoalExplorer reaching only 9.69%, APE 12.82%, and Guardian 17.12% of the predefined target methods across three runs. These findings highlight the need for a hybrid strategy that integrates dynamic execution with path-sensitive static guidance.
4.3. RQ2: Reachability with static analysis
@@
We evaluated the effectiveness of static analysis tools in identifying paths to target methods using FlowDroid (flowdroid) and DroidReach (droidreach). FlowDroid is a well-established static analyzer that is actively maintained and was recently recognized as the best-performing tool for call graph generation (cg_soundness). DroidReach, on the other hand, is the only available tool that reconstructs paths by leveraging call graphs generated via AndroGuard (androguard_cg), applying graph-search algorithms to assess method reachability. In our setup, we improved DroidReach by restricting entry points to only exported components, whereas the original configuration considers all manifest-declared components as potential entry points. We note that DroidReach also supports FlowDroid-generated call graphs, but relies on an outdated version; thus, we excluded that option and ran the tools independently for comparison. We used FlowDroid in its CHA configuration, following a recent comparison study (cg_soundness). We also intended to include GoalExplorer in our static analysis comparison. However, GoalExplorer’s output is limited to the STG, which flags target activities, but does not report how many target APIs were actually found. In addition, the tool currently suffers from broken Maven dependencies, which prevented us from customizing it to produce the statistics required for a fair comparison. As a result, we had to rely on the pre-built executable released in 2024, which restricted our ability to adapt or extend the tool for our evaluation.
@@
Experiments were conducted using the AndroTest benchmark. For each application, 50 randomly selected target methods were used. After execution, we examined the resulting call graphs and computed whether the target methods were reachable. Our results show that both tools face limitations. FlowDroid was able to reach only 58.81% of the target methods, while DroidReach performed worse, reaching just 9.48%. In terms of performance, FlowDroid had an average analysis time of 35.06 seconds, compared to 23.46 seconds for DroidReach.
@@
Importantly, neither tool handles conditional logic along paths, meaning that reachability is assessed without validating whether the conditions required to execute those paths can actually be satisfied. Furthermore, the entry points used by both tools are either artificial or overly broad. More specifically, FlowDroid constructs a synthetic dummyMain method to simulate execution, but it provides no instructions for how to drive it or interpret paths from it. DroidReach, via AndroGuard, restricts itself to exported components declared in the manifest. Since Android applications can have entry points that are programmatically defined, this simplification might lead to missing relevant execution paths entirely.
@@
RQ2: Reachability with static analysis Static analysis tools fail to identify a substantial portion of reachable target methods. FlowDroid reaches only 58.81%, and DroidReach performs even worse by reaching 9.48% of target methods. These findings show that current call graph generation tools are not effective for target-method reachability and motivate the need for more focused and goal-driven approaches.
@@ 5. Design
In this section, we first provide an illustrative example that we will use to explain how GAPS works. Specifically, we first describe an illustrative example to describe how GAPS’ static analysis pre-processing works and how its output guides the automatic interaction.
@@ 5.1. Illustrative Example
Here, we provide an example that will help describe each GAPS component. Furthermore, even with a simple example, it will be possible to immediately grasp the complexity of generating guidance for automatic interaction through static analysis. In our example, we want to build a path to reach the target_method in ActivityA from one of the app entry points, e.g., the MainActivity with its lifecycle. In 1, the MainActivity triggers the display of ActivityA (6 - 9) when the corresponding button is clicked (6). The button click triggers the creation of an Intent message (7) in the callback and dispatches it through the startActivity API (8). Furthermore, a BroadcastReceiver component is registered (5) and can be contacted by sending an Intent message by specifying the com.example.SET_ACCESS action.
@@ 2 Button buttonA = findViewById(button_a);
Finally, in 3, the AccessReceiver class directly influences the reachability of the target method. Upon being contacted, it sets the access variable to true (3), thus affecting the required conditional statement.
@@ 5.2. Overview
Next, GAPS builds feasible paths between entry points and the target method using a lightweight, context-sensitive analysis. For each path, it performs backward slicing from the target method and tracks relevant control-flow transitions. More specifically, a partial call graph is built by concatenating method invocations, involving both explicit and implicit invocations. More details are provided in subsection 5.4. In this example, it reconstructs the sequence from target_method(), which is nested in a conditional branch and the click of the buttonA graphical element. Going further, ActivityA.onStart is only executed through the startActivity(Intent), further nested in the buttonMain click in MainActivity.onStart.
When a conditional statement is encountered, as in line 7 of 2, where access must be true, GAPS analyzes what influences its satisfiability. By performing data-flow analysis through a combination of constant propagation and points-to-analysis, it discovers that the static field ActivityA.access is set to true inside the onReceive method of AccessReceiver. Crucially, this receiver is registered in MainActivity.onStart, and it can be invoked externally by sending a broadcast intent. This process is further described in subsection 5.5. To ensure that the conditional check in ActivityA evaluates to true at runtime, GAPS includes a trigger for contacting the AccessReceiver as part of the execution plan. Furthermore, GUI analysis detects that both buttonMain and buttonA are connected to relevant method calls via click listeners, and therefore must be interacted with during execution. GAPS performs data-flow analysis to retrieve the GUI widgets’ identifiers. Specifically, it analyzes how the widget objects are initialized in the code, tracing their definitions to locate calls to findViewById. In our example, buttonMain and buttonA are both assigned through such calls, and the argument to each function (i.e., button_main and button_a) is recorded as the identifier of the corresponding UI element. Design choices around this process are explained in subsection 5.6.
@@ 5.3. Mapping Inter-Component Communication
GAPS employs a strategy to map ICC based on the understanding that sending an Intent object implies passing it as an argument to a specific Android API call. The primary objective is to extract information regarding both the sender and receiver of the message exchanged through ICC. This way, components that are reachable can be used as app entry points during dynamic interaction, either explicitly (via a fully qualified class name) or implicitly (through an action field).
@@
In the illustrative example, the MainActivity uses the registerReceiver method to instantiate a broadcast receiver, that will accept intent with the “com.example.SET_ACCESS” action, as defined by the corresponding IntentFilter. Additionally, the startActivity API allows to start the ActivityA.

### v2

@@ 3.3. Call Graph Generation
FlowDroid (Arzt et al., 2014) is a taint analysis tool that generates the app’s call graph to discover privacy and security leaks between defined sources and sinks. FlowDroid constructs a call graph using Soot (Soot, 2025b) and performs a flow- and context-sensitive analysis. However, FlowDroid did not model ICC in its initial version, leaving an essential contribution to call graphs unaddressed. This was later introduced with the integration of IccTA (Li et al., 2015). Similarly, Amandroid (Wei et al., 2018) is a taint analysis tool that builds an inter-procedural call graph and computes points-to information for all objects of an app component. FlowDroid and Amandroid employ data-flow analysis techniques to construct a call graph using a forward-visit approach, where every branch is visited during the process. In addition, this approach has proven to be resource-intensive when integrated with the increasing complexity of apps (Avdiienko et al., 2015). GAPS aims for a more scalable approach: we build a partial inter-procedural call graph backward, starting from a target method invocation, to limit processing overhead. DroidReach (Borzacchiello et al., 2022) is the only known path reconstruction tool. Its path reconstruction techniques rely on generating a complete call graph using either the Androguard call graph creation tool (Androguard, 2025c) or FlowDroid.
In subsection 4.3, we compare GAPS against DroidReach and FlowDroid. In both cases, GAPS’s static analysis module proved to be faster and more effective in generating paths towards target methods compared to its counterparts. Additionally, it provides specific app entry points, execution paths to satisfy conditional statements, and corresponding GUI events to guide automated interaction.
@@ 3.4. Hybrid analysis
Andrubis (Lindorfer et al., 2014) is a hybrid analysis tool created for detecting Android malware. Andrubis’ static analysis retrieves meta information (e.g., permissions, components, package name) and examines the Dalvik bytecode to extract a complete list of Java objects and methods. Subsequently, dynamic analysis is aided by information previously collected to interact with known entry points and measure code coverage, but it is not ultimately guided by the data collected during static analysis. AndroPyTool (Martín et al., 2018) is another malware detection tool that combines static and dynamic analysis. It utilizes Androguard to extract essential features (e.g., package name, permissions, opcodes, API calls), and FlowDroid to perform taint analysis. Samples are run inside a sandbox. However, dynamic analysis is not assisted by static analysis; instead, they are included as separate contributions in the output. GAPS aims to use static analysis information to provide path-sensitive instructions and guide automatic interaction. Evodroid (Mahmood et al., 2014) proposes an evolutionary approach to the system testing of Android apps. It builds a Control Graph Model consisting of code segments used during GUI exercising to increase coverage. GAPS, instead, proposes a targeted approach to reach a specific method inside the app. Furthermore, Evodroid bases its static analysis on MoDisco (Eclipse, 2025), which requires the source code of the app to be provided; however, this is not always possible for most apps on the Google Play Store. GAPS can analyze any APK and DEX files as input and does not require the source code to operate. LetterBomb (Garcia et al., 2017) employs a hybrid approach for automatically generating exploits for Android apps, specifically targeting Inter-Component Communication (ICC) vulnerabilities. The methodology combines a path-sensitive symbolic execution-based static analysis with software instrumentation and test oracles to identify and exploit weaknesses in the ICC interface. GAPS differs from LetterBomb in that it performs static analysis without relying on symbolic execution, which is known to be computationally expensive and often impractical for large, real-world apps. Instead, GAPS uses a lightweight static approach to reconstruct paths to a target method and then translates them into high-level instructions to guide UI exploration.
@@
GoalExplorer (Lai and Rubin, 2019) is a hybrid system for automated, goal-driven exploration of Android applications. Its core idea is to first statically model the app’s UI screens and the transitions between them by combining results from modified versions of FlowDroid (Arzt et al., 2014) and BackStage (Kuznetsov et al., 2018), creating an STG. This STG is then used to guide dynamic exploration, allowing the tool to navigate to a specific target of interest, such as an Android activity, an API call, or a program statement. In subsection 4.2, we compared GoalExplorer against GAPS, and it only achieved 9.69% of the target methods on average across three distinct runs.
@@ 4. Motivation
  • RQ1: How effective are GUI testers at triggering the execution of target methods?
  • RQ2: How effective is static analysis in reconstructing a path towards a method?
  • RQ3: How does GAPS perform compared to state-of-the-art approaches?
  • RQ4: Can GAPS be applied to real-world apps?
@@ 4.1. Datasets & Setup
For the first three research questions (RQ1–RQ3), we used 56 applications from the AndroTest benchmark, for which we collected the apps’ latest version. These apps were instrumented using AndroLogs (Samhi and Zeller, 2024) to measure method reachability by logging every time a method was executed. For each app, we randomly selected a consistent set of 50 target methods to be reached during testing, out of which only 34.39% are located inside an Activity.
@@
Static analysis experiments are performed on a CloudLab (Duplyakin et al., 2019) Ubuntu 22.04 machine with 64 GB of RAM, while dynamic ones are performed on an x86-64 emulator running Android 13 and Pixel 2 with Android 11 when an ARM architecture is required.
4.2. RQ1: GUI testers’ effectiveness for dynamically executing target methods
@@
To evaluate the effectiveness of GUI testers in dynamically executing targeted methods within Android applications, we selected APE, recognized as one of the best-performing tools in recent studies (Wang et al., 2021; Su et al., 2021; Akinotcho et al., 2025), and Guardian, an LLM-based UI exploration tool that has outperformed other task-guided approaches (Ran et al., 2024). We also compared against GoalExplorer (Lai and Rubin, 2019), a hybrid analysis tool that statically builds an STG to guide the Stoat (Su et al., 2017a) stochastic GUI testing. The experiment was conducted using the AndroTest benchmark. Each app was executed for five minutes across three independent runs. We chose a 5-minute timeout for all competing tools, including GAPS, to ensure an efficient dynamic analysis. Previous works (Samhi et al., 2025; Ruggia et al., 2021) chose the same threshold. During execution, logs were collected to verify the dynamic reachability of the target methods. To execute Guardian, we tasked the LLM model to “interact with the application to maximize exploration”.
@@
Considering the set of 50 target methods for each app, over three runs, GoalExplorer executed an average of 9.69% methods, APE 12.82%, and Guardian 17.12%. These results demonstrate the differences in usability among the three techniques: APE can rapidly generate GUI events, but struggles to consistently build app states that facilitate the discovery of more hidden parts of the app code. GoalExplorer uses the STG to guide Stoat by flagging activities when a target API is invoked. However, it faces limitations when the target methods reside in third-party libraries or when control-flow conditions are involved, as these often require more complex interactions beyond merely reaching the target activity. Finally, Guardian leverages the LLM’s decision-making capabilities to systematically interact with UI elements in a goal-oriented manner, enabling it to uncover more comprehensive execution paths. Despite this, its average capability to execute target methods remains low, highlighting that semantic planning alone is insufficient without deeper control-flow awareness.
@@
RQ1: GUI testers effectiveness for dynamically executing target methods Our evaluation highlights the limitations of dynamic interaction tools in executing target methods. Both APE and Guardian show low success rates, with GoalExplorer reaching only 9.69%, APE 12.82%, and Guardian 17.12% of the predefined target methods across three runs. These findings highlight the need for a hybrid strategy that integrates dynamic execution with path-sensitive static guidance.
4.3. RQ2: Path reconstruction through static analysis
@@
We evaluated the effectiveness of static analysis tools in identifying paths to target methods using FlowDroid (Arzt et al., 2014) and DroidReach (Borzacchiello et al., 2022). FlowDroid is a well-established static analyzer that is actively maintained and was recently recognized as the best-performing tool for call graph generation (Samhi et al., 2024). DroidReach, on the other hand, is the only available tool that reconstructs paths by leveraging call graphs generated via AndroGuard (Androguard, 2025c), applying graph-search algorithms to assess method reachability. In our setup, we improved DroidReach by restricting entry points to only exported components, whereas the original configuration considers all manifest-declared components as potential entry points. We note that DroidReach also supports FlowDroid-generated call graphs, but relies on an outdated version; thus, we excluded that option and ran the tools independently for comparison. We used FlowDroid in its Class Hierarchy Analysis (CHA) configuration, following a recent comparison study (Samhi et al., 2024), which demonstrated that this setting yields superior results compared to other strategies and tools. We also intended to include GoalExplorer in our static analysis comparison. However, GoalExplorer’s output is limited to the STG, which flags target activities but does not report the number of target APIs that were actually found. In addition, the tool currently suffers from broken Maven dependencies, which prevented us from customizing it to produce the statistics required for a fair comparison. As a result, we had to rely on the pre-built executable released in 2024, which restricted our ability to adapt or extend the tool for our evaluation.
@@
Experiments were conducted using the AndroTest benchmark. For each application, 50 randomly selected target methods were used. After execution, we examined the resulting call graphs and computed whether the target methods were included in the static paths. Our results show that both tools face limitations. FlowDroid was able to generate paths towards only 58.81% of the target methods, while DroidReach performed worse, i.e., just 9.48%. In terms of performance, FlowDroid had an average analysis time of 35.06 seconds, compared to 23.46 seconds for DroidReach.
@@
Importantly, neither tool handles conditional logic along paths, meaning that reachability is assessed without validating whether the conditions required to execute those paths can actually be satisfied. Furthermore, the entry points used by both tools are either artificial or overly broad. More specifically, FlowDroid constructs a synthetic dummyMain method to simulate execution, but it provides no instructions for how to drive it or interpret paths from it. DroidReach, via AndroGuard, restricts itself to exported components declared in the manifest. Since Android applications can have programmatically defined entry points, this simplification may lead to missing relevant execution paths entirely.
@@
RQ2: Path reconstruction through static analysis Static analysis tools fail to build paths towards a substantial portion of the target methods: FlowDroid’s generated paths reach only 58.81% of the methods, while DroidReach only reaches the 9.48%. These findings demonstrate that current call graph generation tools are ineffective for target-method reachability, underscoring the need for more focused and goal-driven approaches.
@@ 5. Design
In this section, we provide an illustrative example that will be used to explain how GAPS works. Specifically, we first describe an illustrative example to describe how GAPS’ static analysis pre-processing works and how its output guides the automatic interaction.
@@ 5.1. Illustrative Example
Here, we provide an example that will help describe each GAPS component. Furthermore, even with a simple example, it will be possible to immediately grasp the complexity of generating guidance for automatic interaction through static analysis. In our example, we aim to create a path to reach the target_method in ActivityA from one of the app’s entry points, such as the MainActivity with its lifecycle. In 1, the MainActivity triggers the display of ActivityA (6 - 9) when the corresponding button is clicked (6). The button click triggers the creation of an Intent message (7) in the callback and dispatches it through the startActivity API (8). Furthermore, a BroadcastReceiver component is registered (5) and can be contacted by sending an Intent message by specifying the com.example.SET_ACCESS action.
@@ 2 Button buttonA = findViewById(button_a);
Finally, in 3, the AccessReceiver class directly controls the invocation of the target method. Upon being contacted, it sets the access variable to true (3), thus affecting the required conditional statement.
@@ 5.2. Overview
Next, GAPS builds feasible paths between entry points and the target method using a lightweight, context-sensitive analysis. For each path, it performs backward slicing from the target method and tracks relevant control-flow transitions. More specifically, a partial call graph is built by concatenating method invocations, involving both explicit and implicit invocations. More details are provided in subsection 5.4. In this example, it reconstructs the sequence from target_method(), which is nested within a conditional branch and triggered by the click of the buttonA graphical element. Going further, ActivityA.onStart is only executed through the startActivity(Intent), further nested in the buttonMain click in MainActivity.onStart.
When a conditional statement is encountered, as in line 7 of 2, where access must be true, GAPS analyzes what influences its satisfiability. By performing data-flow analysis through a combination of constant propagation and points-to-analysis, it discovers that the static field ActivityA.access is set to true inside the onReceive method of AccessReceiver. Crucially, this receiver is registered in MainActivity.onStart, and it can be invoked externally by sending a broadcast intent. This process is further described in subsection 5.5. To ensure that the conditional check in ActivityA evaluates to true at runtime, GAPS includes a trigger for contacting the AccessReceiver as part of the execution plan. Furthermore, GUI analysis detects that both buttonMain and buttonA are connected to relevant method calls via click listeners, and therefore must be interacted with during execution. GAPS performs data-flow analysis to retrieve the identifiers of GUI widgets. Specifically, it analyzes how the widget objects are initialized in the code, tracing their definitions to locate calls to findViewById. In our example, buttonMain and buttonA are both assigned through such calls, and the argument to each function (i.e., button_main and button_a) is recorded as the identifier of the corresponding UI element. Design choices around this process are explained in subsection 5.6.
@@ 5.3. Mapping Inter-Component Communication
GAPS employs a strategy to map ICC based on the understanding that sending an Intent object implies passing it as an argument to a specific Android API call. The primary objective is to extract information regarding both the sender and receiver of the message exchanged through ICC. This way, exported components can be used as app entry points during dynamic interaction, either explicitly (via a fully qualified class name) or implicitly (through an action field).
@@
In the illustrative example, the MainActivity uses the registerReceiver method to instantiate a broadcast receiver that will accept an intent with the “com.example.SET_ACCESS” action, as defined by the corresponding IntentFilter. Additionally, the startActivity API allows starting the ActivityA.